OpenStudy (anonymous):

Suppose a particle is at position 3 on a straight line. The Velocity is given by the equation v=2t+1 where velocity is mph and t is hours. The particle beings to move for 3 hours 1) How many miles did the particle move? 2) Draw the geometry for the problem 3) What position is the particle now?

OpenStudy (abb0t):

If you're given velocity, to find position, you need to integrate this function given to find position at \(t\)

OpenStudy (anonymous):

Alright , but is there any way you can further explain it , it would be really helpful for me~

OpenStudy (anonymous):

integrating the function we get \[s(t)= t^2+t+3\] to find the distance you should find where does the object moves forward and backward by evaluating v(t)=0 and will see that for positive (real) t the object moves forward so we say that d= s(3)-s(0)=3^2+3+1-1= 12miles

OpenStudy (anonymous):

I think the position of the particle is just evaluate s(3)= 15

OpenStudy (anonymous):

I am sorry about the distance, it's 12 but instead of 1-1 make it 3-3